ANALYSIS OF THE ROTATI0NAL STRUCTURE IN SEVERAL BANDS IN THE HIGH-RESOLUTION INFRARED SPECTRA OF BUTADIENE-TRANS,TRANS-1,4-d_{2} AND -CIS,CIS-1,4-d_{2}.

\underline{NORMAN C. CRAIG}, KEITH A. HANSON, RICHARD W. PIERCE, and SCOTT D. SAYLOR, Department of Chemistry, Oberlin College, Oberlin, OH 44074; ROBERT L. SAMS, EMSL, Pacific Northwest National Laboratory, Richland, WA 99352.

Butadiene-trans,trans-1,4-d_{2} and -cis,cis-1,4-d_{2} were prepared from trans,trans-1,4-dichlorobutadiene. Gas-phase spectra were recorded at 0.002 cm^{−1} resolution in the mid-infrared region on a Bruker IFS 120HR FT spectrometer at PNNL. For the trans,trans isotopomer the C-type band at 1002.553 cm^{−1}, which was not perturbed near the band center, was more fully analyzed than another C-type band at 825.5 cm^{−1}, which was extensively perturbed. Ground state rotational constants for this near-symmetric top (κ = -0.9810) were fit to 1828 ground state combination differences (GSCDs) derived from the two bands. A, B and C rotational constants for the ground state are 1.3397796(7), 0.1306985(2), and 0.1191086(3) cm^{−1}, respectively. For the cis,cis isotopomer the C-type band at 808.367 cm^{−1} was analyzed. Fitting 1559 GSCDs gave κ = -0.9707 and 1.1398737(6), 0.1380156(3), and 0.1231302(3) cm^{−1} for A, B and C, respectively. These rotational constants along with those of other deuterium and ^{13}C isotopomers will be used to determine a complete equilibrium rotational structure of butadiene.